OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths place if necessary. 11x2 - x - 3 = 0 A. {-1, 1.2} B. {-11, 13.2} C. {-0.5, 0.6} D. no real solution

OpenStudy (anonymous):

c

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{{1 \pm \sqrt {1 - 4(11 \cdot - 3)} }}{{2 \cdot 11}}\\ = \frac{{1 \pm \sqrt {133} }}{{22}}\\ = 0.57, - 0.48\\ = 0.6, - 0.5 \end{array}\]